OpenStudy (anonymous):

Consider u(x,y) = (x+1)^2 - y^2,show tha u is harmonic

OpenStudy (anonymous):

to show U is harmonic, get so satisfy Uxx + Uyy= 0 Ux= 2(x+1)(1)- 0 = 2x+2, Uxx= 2 Uy=0-2y, Uyy= -2 Uxx + Uyy= 2-2= 0, so U is harmonic

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

You're welcome!